OpenStudy (anonymous):

locate all the critical points of s(t)=(t-1)^2(t+6)^3

OpenStudy (australopithecus):

I assume you are suppose to use calculus?

OpenStudy (anonymous):

yes

OpenStudy (australopithecus):

To acquire the critical points take the derivative set y=0, solve for x, If any of the values of x are not in the domain of the original function ignore them Then, check the domain of the derivative if you find numbers that aren't in the domain of the derivative but are in the domain of the original function they are critical numbers. If any of the values of x are not in the domain of the original function ignore them

OpenStudy (australopithecus):

Do that and I will show you the next step

OpenStudy (australopithecus):

well there isn't really a next step but I will show you how to draw the function from the critical numbers

OpenStudy (anonymous):

wait im confused can you show me all the steps? i dont understand

OpenStudy (australopithecus):

well take the derivative first

OpenStudy (anonymous):

im really bad... 2(t-1)^1*2(t+6)^3+6(t+6)*(t-1)? is thar the start of it?

OpenStudy (australopithecus):

do you want me to show you the easy way to do derivatives?

OpenStudy (australopithecus):

I will just show you my method it is fail proof

OpenStudy (anonymous):

sure

OpenStudy (australopithecus):

You better be grateful haha, this will take a little bit of time

OpenStudy (anonymous):

:) thanks

OpenStudy (australopithecus):

s(t)=(t-1)^2(t+6)^3 so first split the function into smaller functions and take the derivative so we are left with, g(t) = (t-1)^2 and q(t) = (t+6)^3 FIRST PART g(t) = (t-1)^2 We need to use chain rule to figure out what g'(t) is so we simply split this into smaller functions m(t) = t^(2) m'(t) = 2t l(t) = t-1 l'(t) = 1 Now we apply chain rule which is simply (MEMORIZE THIS RULE) g'(t) = m't( l(t) )*l't so after subbing in according to this rule we are left with g(t) = (t-1)^2 g'(t) = 2(t-1)*1 Now we are half way there PART B q(t) = (t+6)^3 same deal, split them u(t) = t^3 u'(t) = 3t^2 o(t) = t + 6 o'(t) = 1 Now chain rule again, q'(t) = u'(o(t))*o'(t) q'(t) = 3(t+6)^2 * 1 so we have q(t) = (t+6)^3 q'(t) = 3(t+6)^2 * 1 Final Part, putting it together g(t) = (t-1)^2 g'(t) = 2(t-1) q(t) = (t+6)^3 q'(t) = 3(t+6)^2 we clearly need to use product rule here so simply just use the rule s'(t) = g'(t)q(t) + q'(t)g(t) s'(t) = 2(t-1)(t+6)^3 + 3(t+6)^2(t-1)^2 so we are left with s(t)=(t-1)^2(t+6)^3 s'(t) = 2(t-1)(t+6)^3 + 3(t+6)^2(t-1)^2

OpenStudy (australopithecus):

It looks long and time consuming but this is the best way to do derivatives, just keep doing them like this it will get very systematic and eventually you will be able to do them in your head really fast

OpenStudy (australopithecus):

Using this method you have no excuses for making a mistake when taking derivatives ha

OpenStudy (anonymous):

wow thats long lol

OpenStudy (australopithecus):

I made a mistake g'(t) = m't( l(t) )*l't should be, g'(t) = m'( l(t) )*l'(t)

OpenStudy (australopithecus):

Yes it is long but you need to see all the steps so you understand the mental process

OpenStudy (anonymous):

yea im saving it lol

OpenStudy (anonymous):

how do we find the critical points now?

OpenStudy (australopithecus):

Finding critical points Step 1 Check domain of s'(t) and s(t) (re-read what I wrote about this above) Both are R so don't worry about it Step 2 (Good to write this on your test or assignment!!!) set s'(t) = 0 so, 0 = 2(t-1)(t+6)^3 + 3(t+6)^2(t-1)^2 solve for t, what ever you get as t are your critical points

OpenStudy (australopithecus):

But yeah, I cant stress this enough do derivatives the long way until you are super comfortable with them, even then it doesn't hurt to do them the long way. Once you do it this way you will never forget how to do them, as there are only like 3 formulas you need to remember.

OpenStudy (anonymous):

can you help me solve it, im getting mixed up with the numbers and exponents

OpenStudy (australopithecus):

Yeah this looks painful to expand

OpenStudy (australopithecus):

Are you in intro to calculus?

OpenStudy (australopithecus):

or is your prof just evil

OpenStudy (anonymous):

i just have an evil professor lol that doesnt explain anything